OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (–2, 0) and (2, 8)? a. y = 2(x + 2) b.y = 2(x – 2) c. y = –2(x – 2) d.y = –2(x + 2)

OpenStudy (anonymous):

First find the slope: y2-y1 / x2-x1 8 - 0 / 2 -(-2) 8/4 2/1 so slope/m = 2 Point slope form: y - y1 = m(x-1) y - 0 = 2(x-1) y = 2(x - 1)

OpenStudy (anonymous):

i dont get it is not in one of the answers

OpenStudy (anonymous):

Whoops sorry! I meant y - y1 = m(x -(-2)) y - 0 = 2(x+2) y = 2(x + 2)

OpenStudy (amorfide):

the slope means the gradient the equation for the gradient is  m=gradient y2=second y coordinate y1=first y coordinate x2=second x coordinate x1=first x coordinate sub in

OpenStudy (amorfide):

equation of a line has an equation  y1= y coordinate x1 = x coordinate (you can use any set of coordinates just make sure the coordinate you use y from you also use x)